OpenStudy (anonymous):

Please help! Solve for the missing length and the other two angles in the triangle below

OpenStudy (anonymous):



OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

First, use the Law of Cosines to find the length of the side opposite to the 5 degree angle.

OpenStudy (anonymous):

it is 15 degrees

OpenStudy (ranga):

okay use 15 degrees in the formula.

OpenStudy (anonymous):

I don't know where a , b , c are though

OpenStudy (ranga):



OpenStudy (anonymous):

ok thanks hold on

OpenStudy (anonymous):

so a^2=3^2 + 4^2 -2(3)(4)cos(15)

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so do I do a=sqrt of all that?

OpenStudy (ranga):

Yes.

OpenStudy (ranga):

After you find 'a', use the law of sines to find angle B: a / sin(A) = b / sin(B) Then angle C = 180 - (angle A + angle B)

OpenStudy (anonymous):

a= 1.348?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

ok, 15.642=sin(B) ?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Solve for angle B.

OpenStudy (anonymous):

I don''t know how @ranga sin^-1(15.642) is undefined

OpenStudy (anonymous):

15.64 was degrees? you want me to do sin^-1(.273)?

OpenStudy (anonymous):

the problem is that sin^-1(15.642) is not showing any number

OpenStudy (ranga):

Oh, Your calculation is wrong. It cannot be 15.642. Redo that calculation because sine cannot be more than 1.

OpenStudy (ranga):

a / sin(A) = b / sin(B) sin(B) = b * sin(A) / a = 3 * sin(15) / 1.348 =

OpenStudy (ranga):

Since angle A is 15 degrees, set your calculator to degree mode.

OpenStudy (anonymous):

31.17 degrees. ok thanks man

OpenStudy (ranga):

I am getting 35.17 degrees.

OpenStudy (anonymous):

sin^-1(.576) = .6138 radians = 31.17 degrees

OpenStudy (ranga):

Check my comment above. The angle 15 is in degrees. SO you have to set the calculator to degree mode BEFORE you calculate 3 * sin(15) / 1.348 Then when you take the inverse the answer will be in degrees because your calculator is already set to degree mode.

OpenStudy (anonymous):

I'm using wolfram I don't have a calculator near me

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

If I use wolfram I still get 35.17 degrees. 3 * sin(15) / 1.348 = 0.576 arcsin(0.576) = 35.17 degrees.